Tickets for a play at the community theater cost $16 for an adult and $2 for a child. If 210 ticktets were sold and the total receipts were $2380, how many of each type of ticket were sold?
Answer by checkley75(3666)

A+C=210, A=210-C
16A+2C=2380 NOW SUBSTITUTE (210-C) FOR A IN THE SECOND EQUATION & SOLVE FOR C THUS:
16(210-C)+2C=2380
3360-16C+2C=2380
-14C=2380-3360
-14C=-980
C=-980/-14
C=70 CHILDREN TICKETS WERE SOLD.
210-70=140 ADULT TICKETS WERE SOLD.
PROOF
16*140=70*2=2380
2240+140=2380
2380=2380